OpenStudy (anonymous):

1. A 12-foot ladder is leaning against a wall that makes a 90-degree angle with the ground. A troublemaker is slowly pulling the bottom of the ladder along the ground, away from the wall at a rate of 2 feet per minute when the top of the ladder is 6 feet above the ground. b. At that moment, what is the rate of change of the angle between the ladder and the ground?

OpenStudy (anonymous):

You need to use the derivative of the Pythagorean theorem. a^2+b^2=c^2 2a da + 2b db = 0 C is a constant, so it goes to 0 db or change along the ground is 2 ft/min 2(6) (da) + 2(b) (2) =0 I then need to find b 6^2+b^2=12^2 b=10.39 ft so da = -4(10.39)/12 da=-3.46 ft/min I think thats what you were looking for with question a

OpenStudy (anonymous):

for b cos(angle)=b/h -sin(angle) dangle=db/h -sin(angle)=1/2 -(1/2) dangle = 2/12 dangle=-1/3